What is the heat flux (in W/m²) through the slab of 10cm thick if temperature drop across the slab is 40°C and the thermal conductivity of slab is 35 W/m.°C.?


1
Expert's answer
2021-08-23T04:49:10-0400

"\phi _q=-k\frac{dT(x)}{dx}\\
\phi _q=-35*\frac{40}{\frac{10}{100}}\\
\phi _q=-14000W/m²\\
\phi _q=14000W/m²"